OpenStudy (anonymous):

what is the derivative of 3^x??

OpenStudy (asnaseer):

\[\begin{align} \text{let }y&=3^x\\ \ln(y)&=\ln(3^x)=x\ln(3)\\ \frac{1}{y}\frac{dy}{dx}&=\ln(3)\\ \frac{dy}{dx}&=y\ln(3)=3^x\ln(3) \end{align}\]

OpenStudy (anonymous):

3^x * ln(x)